OpenStudy (mathmath333):

the series is 1,5,9,13........ which of the following will be the term of this arithmetic progression a 4551 b 10091 c 7881 d 13531 e 13537<---

OpenStudy (cj49):

e is the ans

OpenStudy (mathmath333):

yes 100%

OpenStudy (mathmath333):

what about \[c\] ?

OpenStudy (cj49):

3384*4+1=13537

OpenStudy (cj49):

even C is 1970*4+1

OpenStudy (mathmath333):

so its c or e

ganeshie8 (ganeshie8):

c and e

OpenStudy (cj49):

c wud be 1971 term and e wud be 3385th term

OpenStudy (mathmath333):

how did u calculate 1971 and 3385

OpenStudy (cj49):

a+(n-1)d a=1 n =no of terms d=difference

OpenStudy (cj49):

t(1971)= 1+(1971-1)4 t(1971)=1+1970*4=7880+1=7881

OpenStudy (mathmath333):

tnks

OpenStudy (cj49):

np..